What products are produced by the electrolysis of an aqueous solution of iron (II) nitrate?



1
Expert's answer
2022-05-29T22:35:02-0400


Fe(NO3)2 + H2O —> Fe2+ + 2 NO3 + H2O

1. Cathode:

If the solution contains cations whose standard oxidation potential is smaller than that of hydrogen but greater than that of Al (from Al+ to Pb+), then both hydrogen ions and metal cations are returned to the cathode at the same time:

Fe2+ + 2 e = Fe

2H2O + 2 e = H2 + 2 OH


2. Anode:

If the solution contains oxygen acid anions, the oxidation potential of these anions is greater than the oxidation potential of the water molecule, so the water molecules are first oxidized at the anode:

2 H2O – 4 e = O2 + 4 H+


3. General reaction:

Fe(NO3)2 +2H2O=Fe + H2 + O2 +2H+ +2NO3

ANSWER: cathode: Fe , H2

anode: O2

solution: HNO3